OpenStudy (ny,ny):

if q=6,m=16, what is the value of p?

OpenStudy (ny,ny):

\[4q ^{2}/p=m\]

OpenStudy (lgbasallote):

so substitute \[\implies \frac{4(6)^2}{p} = 16\] what's the problem?

OpenStudy (lgbasallote):

isolating p?

OpenStudy (ny,ny):

4(36)/p=16

OpenStudy (lgbasallote):

right

OpenStudy (ny,ny):

144/p=16 (I think)

OpenStudy (lgbasallote):

correct

OpenStudy (lgbasallote):

so is your problem how to isolate p?

OpenStudy (ny,ny):

yeah

OpenStudy (lgbasallote):

okay \[\frac{4q^2}{p} = m\] multiply both sides by p \[4q^2 = pm\] divide both sides by m \[\frac{4q^2}{m} = p\] does that help?

OpenStudy (ny,ny):

yeahhh... i was thinking it all wrong :P sooo... (doing calculations now...) is p=9?

OpenStudy (ny,ny):

wait....

OpenStudy (ny,ny):

nevermind sorry

OpenStudy (lgbasallote):

yes p = 9 haha

OpenStudy (ny,ny):

Thanks :D yur always so helpful :3

OpenStudy (lgbasallote):

just doing my avatar duties :D

OpenStudy (ny,ny):

Lol :P